# Low-resolution Prior Equilibrium Network for CT Reconstruction

### v1

@@ 3.1 Network architecture
Figure 4 depicts the network structures, which is an end-to-end CT reconstruction method. There are two Resnet blocks in each iteration of the low-resolution prior reconstruction. As shown at the bottom of Figure 4, each block involves a 3-layer network. Then the overall depth of the network is determined by the number of iterations it contains, which is fixed to strike a balance between the receptive fields and the total number of parameters in the network. The inclusion of residual structures [55] is crucial, which can help prevent the loss of fine details, expedite information flow, and enhance the network’s capacity. The residual structures can alleviate training challenges and expedite convergence. We adopt the log-student-t-distribution function as the loss function
$\phi(\bm{x})=\frac{1}{2}\log(1+\bm{x}^{2}),$ (13)
which is differentiable and suitable for modeling the statistics of natural images. As displayed in Figure 4, we set the numbers of channels in each iteration as $6\rightarrow 32\rightarrow 32\rightarrow 5$ for both $R_{\theta}$ and $S_{\theta}$. The convolution operator plays a vital role in feature extraction, which is a set of size $3\times 3$. The convolution stride is set to 1 and the SAME padding strategy is used in implementation. Additionally, we employ the Xavier initialization scheme for the convolution parameters and initialize all biases to zero.
@@ 4 Convergence Analysis
Now we establish conditions that ensure the convergence of the iterates $\bm{x}^{(k+1)}=f_{\theta}\left(\bm{x}^{(k)};\bm{b}\right)$ in Algorithm 1 to a fixed-point $\bm{x}^{(\infty)}$ as the iteration index $k$ approaches infinity. According to the classical fixed-point theory, the convergence of the iterates to a unique fixed-point is guaranteed if the iteration map $f_{\theta}(\cdot;\bm{b})$ is contractive. A contraction condition can be expressed as follows.
Definition 1.
There exists a constant $10\leq c<$ 1 such that $\left|f_{\theta}(\bm{x};\bm{b})-f_{\theta}(\bm{x}^{\prime},\bm{b})\right|\leq c\left|\bm{x}-\bm{x}^{\prime}\right|$ for all $\bm{x}$ and $\bm{x}^{\prime}$.
To ensure the contractivity of the iterative map and the convergence of the fixed-point iterations for Algorithm 1, we establish conditions on the regularization network $R_{\theta}$ and the data item network $S_{\theta}$, respectively. These conditions guarantee that the resulting iteration map satisfies the contractivity condition and, consequently, the fixed-point iterations for these models converge.
In particular, in accordance with [30], we make the assumptions that the regularization network $R_{\theta}$ and the data item network $S_{\theta}$ meet the following condition: there exist positive values $\epsilon_{1}$ and $\epsilon_{2}$ such that for any $\bm{x},\bm{x}^{\prime}\in\mathbb{R}^{N\times N}$, the following inequality holds
$\eqalign{\left\|\left(R_{\theta}-\bm{I}\right)(\bm{x},\bm{D^{\top}x_{l}})-\left(R_{\theta}-\bm{I}\right)\left(\bm{x}^{\prime},\bm{D^{\top}x_{l}}\right)\right\|\leq\epsilon_{1}\left\|\bm{x}-\bm{x}^{\prime}\right\|,\\ \left\|\left(S_{\theta}-\bm{I}\right)(\bm{Ax},\bm{b})-\left(S_{\theta}-\bm{I}\right)\left(\bm{Ax}^{\prime},\bm{b}\right)\right\|\leq\epsilon_{2}\left\|\bm{x}-\bm{x}^{\prime}\right\|.}$ (14)
@@
$\eqalign{\left\|\partial_{x}N_{R_{\theta}}\right\|=\left\|K_{1}^{\top}\partial_{x}\phi_{1}(K_{1}\bm{x})K_{2}^{\top}\partial_{x}\phi_{2}(K_{2}\phi(K_{1}\bm{x}))K_{3}\right\|,\\ \left\|\partial_{x}N_{S_{\theta}}\right\|=\left\|\bm{A}^{\top}K_{1}^{\top}\partial_{x}\phi_{1}(K_{1}\bm{Ax})K_{2}^{\top}\partial_{x}\phi_{2}(K_{2}\phi(K_{1}\bm{Ax}))K_{3}\right\|,}$
where $\left\|\partial_{x}N_{R_{\theta}}\right\|$ and $\left\|\partial_{x}N_{S_{\theta}}\right\|$ can be uniformly bounded independently of $\bm{x}$. Due to $\sup_{\bm{x}}\left|\phi^{\prime}(\bm{x})\right|=0.5$, we have the following convergence result
@@
(Convergence of LRPE with empirical data fidelity). Assume that the observed data is corrupted by additive Gaussian noise, there is $\mathcal{S}=\frac{1}{2}||\bm{Ax}-\bm{b}||_{2}^{2}$ in (9). Let $L=\lambda_{\max}\left(\bm{A}^{\top}\bm{A}\right)$ and $\mu=\lambda_{\min}\left(\bm{A}^{\top}\bm{A}\right)$, where $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ denote the maximum and minimum eigenvalue, respectively. If the step-size parameter $\eta>0$ satisfies $\eta<1/(L+1)$, we can obtain the following inequality
$\left\|f_{\theta}(\bm{x};\bm{b})-f_{\theta}\left(\bm{x}^{\prime};\bm{b}\right)\right\|\leq\underbrace{(1-\eta(1+\mu)+\eta\epsilon)}_{=:\gamma}\left\|\bm{x}-\bm{x}^{\prime}\right\|$
for all $\bm{x},\bm{x}^{\prime}\in\mathbb{R}^{N\times N}$. The coefficient $\gamma$ is less than 1 if $\epsilon_{1}<1+\mu$, in which case the iterates of the LRPE-net converges.
@@
Let $f_{\theta}(\bm{x};\bm{b})$ be the iteration map for the LRPE with empirical data fidelity. The Jacobian of $f_{\theta}(\bm{x};\bm{b})$ with respect to $\bm{x}$ denoted by $\partial_{x}f_{\theta}(\bm{x};\bm{b})$, is given by
$\partial_{x}f_{\theta}(\bm{x};\bm{b})=\left(\bm{I}-\eta\bm{A}^{\top}\bm{A}\right)-\eta\partial_{x}R_{\theta}(\bm{x}),$
@@
$\eqalign{\left\|\partial_{x}f_{\theta}(\bm{x};\bm{b})\right\|&=\left\|\bm{I}-\eta\bm{A}^{\top}\bm{A}-\eta\partial_{x}R_{\theta}(\bm{x})\right\|\\ &=\left\|\bm{I}-\eta\bm{A}^{\top}\bm{A}-\eta\bm{I}+\eta\bm{I}-\eta\partial_{x}R_{\theta}(\bm{x})\right\|\\ &\leq\left\|(1-\eta)\bm{I}-\eta\bm{A}^{\top}\bm{A}\right\|+\eta\left\|\partial_{x}R_{\theta}(\bm{x})-\bm{I}\right\|\\ &\leq\max_{i}\left|(1-\eta)-\eta\lambda_{i}\right|+\eta\epsilon_{1},}$
where $\lambda_{i}$ denotes the $i$ th eigenvalue of $\bm{A}^{\top}\bm{A}$. Here, we use the assumption that the map $\left(R_{\theta}-\bm{I}\right)(\bm{x}):=R_{\theta}(\bm{x})-\bm{x}$ is $\epsilon_{1}$-Lipschitz to guarantee the spectral norm of the Jacobian $\partial_{x}R_{\theta}(\bm{x})-\bm{I}$ to be bounded by $\epsilon_{1}$. By our assumption $\eta<\frac{1}{1+L}$ with $L:=\max_{i}\lambda_{i}$, we have $\eta<\frac{1}{1+\lambda_{i}}$ for all $i$, which implies $(1-\eta)-\eta\lambda_{i}>0$ for all $i$. Therefore, the maximum in (15) is obtained at $\mu:=\min_{i}\lambda_{i}$, which gives
$\left\|\partial_{x}f_{\theta}(\bm{x};\bm{b})\right\|\leq 1-\eta(1+\mu)+\eta\epsilon_{1}.$
@@
$\left\|f_{\theta}(\bm{x};\bm{b})-f_{\theta}\left(\bm{x}^{\prime};\bm{b}\right)\right\|\leq\underbrace{(1-\eta(1+\mu)+\eta(L\epsilon_{2}+\epsilon_{1}))}_{=:\gamma}\left\|\bm{x}-\bm{x}^{\prime}\right\|$
for all $\bm{x},\bm{x}^{\prime}\in\mathbb{R}^{N\times N}$. The coefficient $\gamma$ is less than 1 if $\epsilon_{1}+L\epsilon_{2}<1+\mu$, in which case the iterates of the LRPE-net converges.
@@
Let $f_{\theta}(\bm{x};\bm{b})$ be the iteration map for LRPE with learned data fidelity. The Jacobian of $f_{\theta}(\bm{x};\bm{b})$ with respect to $\bm{x}\in\mathbb{R}^{N\times N}$, denoted by $\partial_{x}f_{\theta}(\bm{x};\bm{b})$, is given by
$\partial_{x}f_{\theta}(\bm{x};\bm{b})=\left(\bm{I}-\eta\bm{A}^{\top}\partial_{x}S_{\theta}(\bm{Ax})\bm{A}\right)-\eta\partial_{x}R_{\theta}(\bm{x}),$
where $\partial_{x}R_{\theta}(\bm{x})$ is the Jacobian of $R_{\theta}$ with respect to $\bm{x}\in\mathbb{R}^{N\times N}$ and $\partial_{x}D_{\theta}(\bm{Ax})$ is the Jacobian of $D_{\theta}$. Similarly, we show $\left\|\partial_{x}f_{\theta}(\bm{x};\bm{b})\right\|<1$ for all $\bm{x}\in\mathbb{R}^{N\times N}$. Since $\left(R_{\theta}-\bm{I}\right)(\bm{x})$ and $\left(S_{\theta}-\bm{I}\right)(\bm{x})$ are $\epsilon_{1}$-Lipschitz and $\epsilon_{2}$-Lipschitz, respectively, we have the spectral norm of its Jacobian $\partial_{x}R_{\theta}(\bm{x})-\bm{I}$ is bounded by $\epsilon_{1}$ and $\partial_{x}S_{\theta}(\bm{x})-\bm{I}$ is bounded by $\epsilon_{2}$. Then there is
$\eqalign{\left\|\partial_{x}f_{\theta}(\bm{x};\bm{b})\right\|&=\left\|\bm{I}-\eta\bm{A}^{\top}\partial_{x}S_{\theta}(\bm{Ax})\bm{A}-\eta\partial_{x}R_{\theta}(x)\right\|\\ &=\left\|\bm{I}-\eta\bm{I}+\eta\bm{I}-\eta\bm{A}^{\top}\bm{A}+\eta\bm{A}^{\top}\bm{A}-\eta\bm{A}^{\top}\partial_{x}S_{\theta}(\bm{Ax})\bm{A}-\eta\partial_{x}R_{\theta}(\bm{x})\right\|\\ &=\left\|\bm{I}-\eta\bm{A}^{\top}\bm{A}-\eta\bm{I}-\eta\left(\partial_{x}R_{\theta}(x)-\bm{I}\right)-\eta\bm{A}^{\top}(\partial_{x}S_{\theta}(\bm{Ax})-\bm{I})\bm{A}\right\|\\ &\leq\left\|(1-\eta)\bm{I}-\eta\bm{A}^{\top}\bm{A}\right\|+\eta\left\|\partial_{x}R_{\theta}(\bm{x})-\bm{I}\right\|+\eta\bm{A}^{\top}\left\|\partial_{x}S_{\theta}(\bm{Ax})-\bm{I}\right\|\bm{A}\\ &\leq\max_{i}\left|(1-\eta)-\eta\lambda_{i}\right|+\eta\epsilon_{1}+\eta L\epsilon_{2},}$
@@
$\left\|\partial_{x}f_{\theta}(\bm{x};\bm{b})\right\|\leq 1-\eta(1+\mu)+\eta\epsilon_{1}+\eta L\epsilon_{2}.$
@@ 5 Numerical Results
In this section, we evaluate our LRPE on both sparse-view and limited-angle reconstruction problems and compare it with several state-of-the-art methods using a dataset of human phantoms. We utilize the peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) as evaluation metrics to measure the quality of the reconstructed images produced by different methods.
@@ 5.1 Comparison algorithms
We employ several recent CT reconstruction methods, including both variational and learning-based approaches, as described below:
@@ 2. $\bullet$
    PD: the Learned Primal-Dual network in [57]. The network is a deep unrolled neural network with 10 iterations. The number of initialization channels for both primal and dual values is set to 5. The network parameters are initialized using the Xavier initialization scheme. In all experiments, we employed the mean squared loss as the objective function, which measures the discrepancy between the reconstructed image and the ground truth.

### v2

@@ 3.1 Network architecture
Figure 5 depicts the network structures, which is an end-to-end CT reconstruction method. There are two Resnet blocks in each stage of the low-resolution prior reconstruction. As shown at the bottom of Figure 5, each block involves a 3-layer network. Then the overall depth of the network is determined by the number of stages it contains, which is fixed to strike a balance between the receptive fields and the total number of parameters in the network. The inclusion of residual structures [30] is crucial, which can help prevent the loss of fine details, expedite information flow, and enhance the network’s capacity. The residual structures can alleviate training challenges and expedite convergence. We adopt the log-student-t-distribution function as the loss function
$\phi({x})=\frac{1}{2}\log(1+{x}^{2}),$ (12)
which is differentiable and suitable for modeling the statistics of natural images. We set the numbers of channels in each stage as $6\rightarrow 32\rightarrow 32\rightarrow 5$ for both $S_{\theta}$ and $R_{\theta}$. The convolution operator plays a vital role in feature extraction, where we set the size of the convolution kernel to 3 in implementation. The convolution stride is set to 1 and the SAME padding strategy is used in implementation. Additionally, we employ the Xavier initialization scheme for the convolution parameters and initialize all biases to zero.
@@ 4 Convergence Analysis
According to the classical fixed-point theorem [23], the convergence of the iterates to a unique fixed-point is guaranteed if the iteration map $f_{\theta}(\cdot;\bm{b})$ is contractive, i.e. if exists a constant $10\leq c<$ 1 such that $\left\|f_{\theta}(\bm{u};\bm{b})-f_{\theta}(\bm{u}^{\prime},\bm{b})\right\|\leq c\left\|\bm{u}-\bm{u}^{\prime}\right\|$ for all $\bm{u}$ and $\bm{u}^{\prime}$. Now we establish conditions that ensure the convergence of the iterates $\bm{u}^{(k+1)}=f_{\theta}\left(\bm{u}^{(k)};\bm{b}\right)$ in Algorithm 1 to a fixed-point $\bm{u}^{(\infty)}$ as the stage index $k$ approaches infinity. More specifically, we establish conditions on the regularization network $R_{\theta}$ and the data item network $S_{\theta}$, respectively. These conditions guarantee that the resulting stage map satisfies the contractivity condition and, consequently, the fixed-point iterations for these models converge.
In accordance with [10], in the absence of noise, we make the assumptions that the regularization network $R_{\theta}$ and the data item network $S_{\theta}$ meet the following conditions: there exist positive values $\epsilon_{1}$ and $\epsilon_{2}$ such that for any $\bm{u},\bm{u}^{\prime}\in\mathbb{R}^{N^{2}}$, the following inequality holds
$\eqalign{\left\|\left(R_{\theta}-\bm{I}\right)(\bm{u},\bm{D^{\top}u_{l}})-\left(R_{\theta}-\bm{I}\right)\left(\bm{u}^{\prime},\bm{D^{\top}u_{l}}\right)\right\|\leq\epsilon_{1}\left\|\bm{u}-\bm{u}^{\prime}\right\|,\\ \left\|\left(S_{\theta}-\bm{I}\right)(\bm{Au},\bm{b})-\left(S_{\theta}-\bm{I}\right)\left(\bm{Au}^{\prime},\bm{b}\right)\right\|\leq\epsilon_{2}\left\|\bm{u}-\bm{u}^{\prime}\right\|.}$ (13)
@@
$\eqalign{\left\|\partial_{u}N_{R_{\theta}}\right\|=\left\|K_{1}^{\top}\partial_{u}\phi_{1}(K_{1}\bm{u})K_{2}^{\top}\partial_{u}\phi_{2}(K_{2}\phi(K_{1}\bm{u}))K_{3}\right\|,\\ \left\|\partial_{u}N_{S_{\theta}}\right\|=\left\|\bm{A}^{\top}K_{1}^{\top}\partial_{u}\phi_{1}(K_{1}\bm{Au})K_{2}^{\top}\partial_{u}\phi_{2}(K_{2}\phi(K_{1}\bm{Au}))K_{3}\right\|,}$
due to $\sup_{\bm{u}}\left|\phi^{\prime}(\bm{u})\right|=0.5$, $\left\|\partial_{u}N_{R_{\theta}}\right\|$ and $\left\|\partial_{u}N_{S_{\theta}}\right\|$ can be uniformly bounded independently of $\bm{u}$. We have the following convergence result
@@
(Convergence of LRPE with empirical data fidelity). Assume that the observed data is without noise, there is $\mathcal{S}=\frac{1}{2}||\bm{Au}-\bm{b}||_{2}^{2}$ in (9). Let $L=\lambda_{\max}\left(\bm{A}^{\top}\bm{A}\right)$ and $\mu=\lambda_{\min}\left(\bm{A}^{\top}\bm{A}\right)$, where $\lambda_{\max}(\cdot)$ and $\lambda_{\min}(\cdot)$ denote the maximum and minimum eigenvalue, respectively. If the step-size parameter $\eta>0$ satisfies $\eta<1/(L+1)$, we can obtain the following inequality
$\left\|f_{\theta}(\bm{u};\bm{b})-f_{\theta}\left(\bm{u}^{\prime};\bm{b}\right)\right\|\leq\underbrace{(1-\eta(1+\mu)+\eta\epsilon_{1})}_{=:\gamma}\left\|\bm{u}-\bm{u}^{\prime}\right\|$
for all $\bm{u},\bm{u}^{\prime}\in\mathbb{R}^{N^{2}}$. The coefficient $\gamma$ is less than 1 if $\epsilon_{1}<1+\mu$, in which case the iterates of the LRPE model converges.
@@
Let $f_{\theta}(\bm{u};\bm{b})$ be the iteration map for the LRPE with empirical data fidelity. The Jacobian of $f_{\theta}(\bm{u};\bm{b})$ with respect to $\bm{u}$ denoted by $\partial_{u}f_{\theta}(\bm{u};\bm{b})$, is given by
$\partial_{u}f_{\theta}(\bm{u};\bm{b})=\left(\bm{I}-\eta\bm{A}^{\top}\bm{A}\right)-\eta\partial_{u}R_{\theta}(\bm{u}),$
@@
$\eqalign{\left\|\partial_{u}f_{\theta}(\bm{u};\bm{b})\right\|&=\left\|\bm{I}-\eta\bm{A}^{\top}\bm{A}-\eta\partial_{u}R_{\theta}(\bm{u})\right\|\\ &=\left\|\bm{I}-\eta\bm{A}^{\top}\bm{A}-\eta\bm{I}+\eta\bm{I}-\eta\partial_{u}R_{\theta}(\bm{u})\right\|\\ &\leq\left\|(1-\eta)\bm{I}-\eta\bm{A}^{\top}\bm{A}\right\|+\eta\left\|\partial_{u}R_{\theta}(\bm{u})-\bm{I}\right\|\\ &\leq\max_{i}\left|(1-\eta)-\eta\lambda_{i}\right|+\eta\epsilon_{1},}$
where $\lambda_{i}$ denotes the $i$ th eigenvalue of $\bm{A}^{\top}\bm{A}$. Here, we use the assumption that the map $\left(R_{\theta}-\bm{I}\right)(\bm{u}):=R_{\theta}(\bm{u})-\bm{u}$ is $\epsilon_{1}$-Lipschitz to guarantee the spectral norm of the Jacobian $\partial_{u}R_{\theta}(\bm{u})-\bm{I}$ to be bounded by $\epsilon_{1}$. By our assumption $\eta<\frac{1}{1+L}$ with $L:=\max_{i}\lambda_{i}$, we have $\eta<\frac{1}{1+\lambda_{i}}$ for all $i$, which implies $(1-\eta)-\eta\lambda_{i}>0$ for all $i$. Therefore, the maximum in (14) is obtained at $\mu:=\min_{i}\lambda_{i}$, which gives
$\left\|\partial_{u}f_{\theta}(\bm{u};\bm{b})\right\|\leq 1-\eta(1+\mu)+\eta\epsilon_{1}.$
@@
$\left\|f_{\theta}(\bm{u};\bm{b})-f_{\theta}\left(\bm{u}^{\prime};\bm{b}\right)\right\|\leq\underbrace{(1-\eta(1+\mu)+\eta(L\epsilon_{2}+\epsilon_{1}))}_{=:\gamma}\left\|\bm{u}-\bm{u}^{\prime}\right\|$
for all $\bm{u},\bm{u}^{\prime}\in\mathbb{R}^{N^{2}}$. The coefficient $\gamma$ is less than 1 if $\epsilon_{1}+L\epsilon_{2}<1+\mu$, in which case the iterates of the LRPE-net converges.
@@
Let $f_{\theta}(\bm{u};\bm{b})$ be the iteration map for LRPE with learned data fidelity. The Jacobian of $f_{\theta}(\bm{u};\bm{b})$ with respect to $\bm{u}\in\mathbb{R}^{N^{2}}$, denoted by $\partial_{u}f_{\theta}(\bm{u};\bm{b})$, is given by
$\partial_{u}f_{\theta}(\bm{u};\bm{b})=\left(\bm{I}-\eta\bm{A}^{\top}\partial_{u}S_{\theta}(\bm{Au})\bm{A}\right)-\eta\partial_{u}R_{\theta}(\bm{u}),$
where $\partial_{u}R_{\theta}(\bm{u})$ is the Jacobian of $R_{\theta}$ with respect to $\bm{u}\in\mathbb{R}^{N^{2}}$ and $\partial_{u}S_{\theta}(\bm{Au})$ is the Jacobian of $S_{\theta}$. Similarly, we show $\left\|\partial_{u}f_{\theta}(\bm{u};\bm{b})\right\|<1$ for all $\bm{u}\in\mathbb{R}^{N^{2}}$. Since $\left(R_{\theta}-\bm{I}\right)(\bm{u})$ and $\left(S_{\theta}-\bm{I}\right)(\bm{u})$ are $\epsilon_{1}$-Lipschitz and $\epsilon_{2}$-Lipschitz, respectively, we have the spectral norm of its Jacobian $\partial_{u}R_{\theta}(\bm{u})-\bm{I}$ is bounded by $\epsilon_{1}$ and $\partial_{u}S_{\theta}(\bm{u})-\bm{I}$ is bounded by $\epsilon_{2}$. Then there is
$\eqalign{\left\|\partial_{u}f_{\theta}(\bm{u};\bm{b})\right\|&=\left\|\bm{I}-\eta\bm{A}^{\top}\partial_{u}S_{\theta}(\bm{Au})\bm{A}-\eta\partial_{u}R_{\theta}(\bm{u})\right\|\\ &=\left\|\bm{I}-\eta\bm{I}+\eta\bm{I}-\eta\bm{A}^{\top}\bm{A}+\eta\bm{A}^{\top}\bm{A}-\eta\bm{A}^{\top}\partial_{u}S_{\theta}(\bm{Au})\bm{A}-\eta\partial_{u}R_{\theta}(\bm{u})\right\|\\ &=\left\|\bm{I}-\eta\bm{A}^{\top}\bm{A}-\eta\bm{I}-\eta\left(\partial_{u}R_{\theta}(u)-\bm{I}\right)-\eta\bm{A}^{\top}(\partial_{u}S_{\theta}(\bm{Au})-\bm{I})\bm{A}\right\|\\ &\leq\left\|(1-\eta)\bm{I}-\eta\bm{A}^{\top}\bm{A}\right\|+\eta\left\|\partial_{u}R_{\theta}(\bm{u})-\bm{I}\right\|+\eta\bm{A}^{\top}\left\|\partial_{u}S_{\theta}(\bm{Au})-\bm{I}\right\|\bm{A}\\ &\leq\max_{i}\|(1-\eta)-\eta\lambda_{i}\|+\eta\epsilon_{1}+\eta L\epsilon_{2},}$
@@
$\left\|\partial_{u}f_{\theta}(\bm{u};\bm{b})\right\|\leq 1-\eta(1+\mu)+\eta\epsilon_{1}+\eta L\epsilon_{2}.$
@@ 5 Numerical Results
In this section, we evaluate our LRPE model on both sparse-view and limited-angle reconstruction problems and compare it with several state-of-the-art methods using a dataset of human phantoms. We utilize the peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) as evaluation metrics to measure the quality of the reconstructed images produced by different methods.
@@ 5.1 Comparison algorithms
We employ several established CT reconstruction methods for comparison studies, including both variational and learning-based approaches, as described below:
@@ 2. $\bullet$
    PD: the Learned Primal-Dual network in [1]. The network is a deep unrolled neural network with 10 stages. The number of initialization channels for both primal and dual values is set to 5. The network parameters are initialized using the Xavier initialization scheme. In all experiments, we employed the mean squared loss as the objective function, which measures the discrepancy between the reconstructed image and the ground truth.